# LayerKV: Optimizing Large Language Model Serving with Layer-wise KV Cache Management

### v1

@@ Abstract.
The expanding context windows in large language models (LLMs) have greatly enhanced their capabilities in various applications, but they also introduce significant challenges in maintaining low latency, particularly in Time to First Token (TTFT). This paper identifies that the sharp rise in TTFT as context length increases is predominantly driven by queuing delays, which are caused by the growing demands for GPU Key-Value (KV) cache allocation clashing with the limited availability of KV cache blocks. To address this issue, we propose LayerKV, a simple yet effective plug-in method that effectively reduces TTFT without requiring additional hardware or compromising output performance, while seamlessly integrating with existing parallelism strategies and scheduling techniques. Specifically, LayerKV introduces layer-wise KV block allocation, management, and offloading for fine-grained control over system memory, coupled with an SLO-aware scheduler to optimize overall Service Level Objectives (SLOs). Comprehensive evaluations on representative models, ranging from 7B to 70B parameters, across various GPU configurations, demonstrate that LayerKV improves TTFT latency up to 11x and reduces SLO violation rates by 28.7%, significantly enhancing the user experience
@@ 1. Introduction
  • We conduct comprehensive evaluations on models ranging from 7B to 70B across single and multiple GPUs, demonstrating significant optimizations. LayerKV achieves up to 11x improvements in TTFT, reducing Service SLO violation rates by 28.7%, thereby significantly enhancing the user experience.
@@ 2.2. Motivation
Firstly, Figure LABEL:fig:subfig1 illustrates the current phase, where the system leverages PagedAttention GPU kernel to handle decoding iterations with non-contiguous stored KV caches. PagedAttention reserves a significant portion of GPU memory for KV blocks, intended to store future KV cache entries. To determine the amount of memory to allocate for KV blocks, the system profiles the available GPU memory during initialization based on the maximum configured input size. During this process, a fixed proportion (e.g., 90%) of the remaining memory—after accounting for model parameters and activations—is reserved for KV blocks. As context window becomes longer, maximum input configurations correspondingly expands, resulting in greater activation memory usage during profiling. Consequently, the GPU memory for KV blocks decreases. This figure deliberately displays a small number of KV blocks to highlight that the capacity of KV blocks can be significantly limited.
Secondly, Figures. LABEL:fig:subfig2 and LABEL:fig:subfig3 illustrate scenarios where requests of varying lengths are queued for scheduling. Existing serving systems are stateless across requests. In other words, they de-allocate all the cache slots used by a request as soon as it finishes. Therefore, within the iterative batch processing approach, the system allows new requests to initiate the prefill stage earlier if sufficient KV blocks are available, thus reducing queuing delays. To determine whether a new request can be inserted, the system compares the total KV blocks required for its prefill stage with the currently available KV blocks. As a result, if requests have shorter prompts, they can be scheduled immediately, as indicated by the green segment. However, if requests have longer prompts, they must wait until KV blocks are released, necessitating at least one request to be fully completed first. This process can be time-consuming, which consequently results in subsequent requests remaining queued.
@@ 5.2.3. Performance Comparison Under Varying Request Arrival Rates
Figure 6 compares the performance of LayerKV and vLLM across different request arrival rates using the ShareGPT dataset. As the request arrival rate increases, vLLM’s TTFT rises sharply, particularly at higher rates, where queuing delays lead to significant latency spikes. In contrast, LayerKV effectively controls TTFT, maintaining relatively low latency even under heavy loads. In this case, LayerKV achieves up to an 11x reduction in TTFT latency.
@@ 6.1. KV Cache Optimization
System-level optimizations focus on increasing available memory capacity. A straightforward approach is to utilize multiple GPUs or offload KV caches to CPU memory or even disk when GPU memory is insufficient. Various parallel strategies partition memory demands across token-wise (Wu et al., 2024; Liu et al., 2023; Brandon et al., 2023), model-wise (Zheng et al., 2022; Narayanan et al., 2019), operator-wise (Shoeybi et al., 2019; Pope et al., 2023), and stage-wise (Qin et al., 2024; Zhong et al., 2024) levels, effectively distributing memory pressure. However, these strategies typically require multi-GPU or distributed environments and are not applicable to single-device setups. On the other hand, well-established offloading methods (Lee et al., 2024; Sheng et al., 2023) are more suitable for offline scenarios, where the primary focus is on optimizing system throughput rather than SLO metrics. In contrast, LayerKV maintains lossless precision, supports both single- and multi-GPU setups, and simultaneously accounts for SLO requirements.
@@ 6.2. LLM Serving Systems
Table 1 compares LayerKV with state-of-the-art LLM serving systems. vLLM proposes a PagedAttention mechanism to minimize GPU memory fragmentation, boosting batch size and token generation throughput. DistServe (Zhong et al., 2024) divides prefill and decode execution to different GPUs in order to avoid performance interference of the two computation stages. DeepSpeed-FastGen (Holmes et al., 2024) and Sarathi (Agrawal et al., 2023) decompose prompts into small chunks and combine them with decode tokens, which improves responsiveness and tail latency.
@@ 7. Conclusion
To address the significant challenge of increasing Time to First Token (TTFT) in LLM serving under large context lengths, we have developed LayerKV, which introduces layer-wise KV block allocation, management, and offloading for fine-grained control over system memory. LayerKV effectively reduces queuing delays by optimizing GPU KV cache usage without requiring additional hardware or compromising output quality. Our comprehensive evaluations on models ranging from 7B to 70B parameters across single and multiple GPUs demonstrate that LayerKV achieves up to an 11x reduction in TTFT latency and reduces Service Level Objective (SLO) violation rates by 28.7%, significantly enhancing user experience.

### v3

@@ Abstract.
The expanding context windows in large language models (LLMs) have greatly enhanced their capabilities in various applications, but they also introduce significant challenges in maintaining low latency, particularly in Time to First Token (TTFT). This paper identifies that the sharp rise in TTFT as context length increases is predominantly driven by queuing delays, which are caused by the growing demands for GPU Key-Value (KV) cache allocation clashing with the limited availability of KV cache blocks. To address this issue, we propose LayerKV, a simple yet effective plug-in method that effectively reduces TTFT without requiring additional hardware or compromising output performance, while seamlessly integrating with existing parallelism strategies and scheduling techniques. Specifically, LayerKV introduces layer-wise KV block allocation, management, and offloading for fine-grained control over system memory, coupled with an SLO-aware scheduler to optimize overall Service Level Objectives (SLOs). Comprehensive evaluations on representative models, ranging from 7B to 70B parameters, across various GPU configurations, demonstrate that LayerKV improves TTFT latency up to 69x and reduces SLO violation rates by 28.7%, significantly enhancing the user experience.
@@ 1. Introduction
  • We conduct comprehensive evaluations on models ranging from 7B to 70B across single and multiple GPUs, demonstrating significant optimizations. LayerKV achieves up to 69x improvements in TTFT, reducing Service SLO violation rates by 28.7%, thereby significantly enhancing the user experience.
@@ 2.2. Motivation
Firstly, Figure 2 (a) illustrates the current phase, where the system leverages PagedAttention GPU kernel to handle decoding iterations with non-contiguous stored KV caches. PagedAttention reserves a significant portion of GPU memory for KV blocks, intended to store future KV cache entries. To determine the amount of memory to allocate for KV blocks, the system profiles the available GPU memory during initialization based on the maximum configured input size. During this process, a fixed proportion (e.g., 90%) of the remaining memory—after accounting for model parameters and activations—is reserved for KV blocks. As context window becomes longer, maximum input configurations correspondingly expands, resulting in greater activation memory usage during profiling. Consequently, the GPU memory for KV blocks decreases. This figure deliberately displays a small number of KV blocks to highlight that the capacity of KV blocks can be significantly limited.
Secondly, Figures 2 (b) and (c) illustrate scenarios where requests of varying lengths are queued for scheduling. Existing serving systems are stateless across requests. In other words, they de-allocate all the cache slots used by a request as soon as it finishes. Therefore, within the iterative batch processing approach, the system allows new requests to initiate the prefill stage earlier if sufficient KV blocks are available, thus reducing queuing delays. To determine whether a new request can be inserted, the system compares the total KV blocks required for its prefill stage with the currently available KV blocks. As a result, if requests have shorter prompts, they can be scheduled immediately, as indicated by the green segment. However, if requests have longer prompts, they must wait until KV blocks are released, necessitating at least one request to be fully completed first. This process can be time-consuming, which consequently results in subsequent requests remaining queued.
@@ 5.2.3. Performance Comparison Under Varying Request Arrival Rates
Figure 7. P99 TTFT Comparison of LayerKV and vLLM Under Varying Request Arrival Rates.
Figure 6 compares the performance of LayerKV and vLLM across different request arrival rates using the ShareGPT dataset. As the request arrival rate increases, vLLM’s TTFT rises sharply, particularly at higher rates, where queuing delays lead to significant latency spikes. In contrast, LayerKV effectively controls TTFT, maintaining relatively low latency even under heavy loads. In this case, LayerKV achieves up to an 69x reduction in mean TTFT latency and 45x reduction in P99 TTFT latency (by Figure 7).
@@ 6.1. KV Cache Optimization
System-level optimizations focus on increasing available memory capacity. A straightforward approach is to utilize multiple GPUs or offload KV caches to CPU memory or even disk when GPU memory is insufficient. Various parallel strategies partition memory demands across token-wise (Wu et al., 2024; Liu et al., 2023; Brandon et al., 2023), model-wise (Zheng et al., 2022; Narayanan et al., 2019), operator-wise (Shoeybi et al., 2019; Pope et al., 2023), and stage-wise (Qin et al., 2024; Zhong et al., 2024) levels, effectively distributing memory pressure. However, these strategies typically require multi-GPU or distributed environments and are not applicable to single-device setups. On the other hand, well-established offloading methods (Lee et al., 2024; Sheng et al., 2023) are more suitable for offline scenarios, where the primary focus is on optimizing system throughput rather than SLO metrics. In contrast, LayerKV maintains lossless precision, supports both single- and multi-GPU setups, and simultaneously accounts for SLO requirements. vTensor (Xu et al., 2024) proposes a new virtual memory abstraction for managing KV cache memory using CUDA virtual memory management, enabling both memory defragmentation and computation flexibility. This optimization is orthogonal to LayerKV.
@@ 6.2. LLM Serving Systems
Table 1 compares LayerKV with state-of-the-art LLM serving systems. vLLM (Kwon et al., 2023) proposes a PagedAttention mechanism to minimize GPU memory fragmentation, boosting batch size and token generation throughput. DistServe (Zhong et al., 2024) divides prefill and decode execution to different GPUs in order to avoid performance interference of the two computation stages. DeepSpeed-FastGen (Holmes et al., 2024) and Sarathi (Agrawal et al., 2023) decompose prompts into small chunks and combine them with decode tokens, which improves responsiveness and tail latency.
@@ 7. Conclusion
To address the significant challenge of increasing Time to First Token (TTFT) in LLM serving under large context lengths, we have developed LayerKV, which introduces layer-wise KV block allocation, management, and offloading for fine-grained control over system memory. LayerKV effectively reduces queuing delays by optimizing GPU KV cache usage without requiring additional hardware or compromising output quality. Our comprehensive evaluations on models ranging from 7B to 70B parameters across single and multiple GPUs demonstrate that LayerKV achieves up to an 69x reduction in TTFT latency and reduces Service Level Objective (SLO) violation rates by 28.7%, significantly enhancing user experience.